Hi . i have a qustion that i don't know how to splve it please help me
İ can`t speak english good but i will try translte it english
there is a a pressure of 100 Pa air in bowl on (0 C) then add 20gram water and raising the temperature until(?)---->100C and what is the last pressure if u understand what i mean help me please
1
Expert's answer
2012-07-19T07:39:27-0400
P1=P/(mT)=100/(0.0020*100)=50 Pa